Two balls, each of radius R, equal mass and density are placed in contact, then the force of gravitation between them is proportional to –?

Grade:

2 Answers

Kaustubh Nayyar
9 years ago
F = ( G m2 ) / (2R) = Gm/ 4R2
Hence propotional to 1/4R2
Mitansh Shroff
7 years ago
F=GM2/(2R)2
  =GM2/4R2
Therefore, M=(density)*Volume
M=(density)*4piR3/3
M2=(density)2.(4pi/3)2.R6
Here,F=G.(density)2.(4pi/3)2.R
Therefore F is proportional to R4
---------------------------------------